# Graphical Scattering Equations

### v1

@@ 3 The Scattering Matroid
We work in the setting of (5), where $\mathcal{V}_{G}\subset\mathcal{K}_{G}\times\mathcal{M}_{0,n}$. Our hypothesis says that the map to $\mathcal{K}_{G}$ is finite-to-one, and in particular it is dominant onto $\mathcal{K}_{G}$. The first condition means that the codimension of the scattering correspondence in $\mathcal{K}_{G}\times\mathcal{M}_{0,n}$ equals $n-3={\rm dim}(\mathcal{M}_{0,n})$. Its codimension in $\mathbb{P}^{|G|-1}\times\mathcal{M}_{0,n}$ is therefore ${\rm codim}(\mathcal{K}_{G})+n-3=\kappa(G)+n-3$. The latter codimension is precisely the rank of the scattering matroid $\mathcal{S}(G)$, thus (17) holds.
@@
A connected graph $G\subseteq\binom{[n]}{2}$ is called matroidally copious if both (17) and (18)=(19) are satisfied, and the rank of the scattering matroid satisfies the upper bound
@@
[…]
@@ 5 Partial Compactifications
A graph $G\subseteq\binom{[n]}{2}$ is topologically copious if the map (30) is injective. This property is independent of the choice of basis $u_{1},\ldots,u_{r}$ for the kinematic lattice $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$.
@@ 6 Graphs with Few Vertices
The list of topologically copious graphs, computed as described above, coincides with the list of copious graphs for $n=6,7,8$. This gives us considerable confidence in the conjecture that condition 4 is equivalent to conditions 1,2,3 in Theorem 1.3 for arbitrary graphs $G$.
@@
$\small\begin{matrix}\hskip 18.49988pt|G|\backslash\mu(G)\!\!\!\!&1&2&3&4&\!5,\!6\!&\!\!7..9\!\!&\!\!10..24\!\!\\ \!\!12&\!10&2&0&0&0&0&0\\ \!\!13&\!14&8&1&0&0&0&0\\ \!\!14&\!11&13&5&2&0&0&0\\ \!\!15&5&9&5&6&3&0&0\\ \!\!16&0&4&2&5&6&0&0\\ \!\!17&0&0&0&3&2&4&0\\ \!\!18..21&0&0&0&0&1&2&6\end{matrix}\qquad\,\,\,\,\footnotesize\begin{matrix}\qquad|G|\backslash\mu(G)\!\!\!\!&1&2&3,\!4&5,\!6&\!\!7..10\!\!&\!\!11..20\!\!&\!\!21..120\!\!\\ 13&0&1&0&0&0&0&0\\ 14&52&29&3&0&0&0&0\\ 15&\!110&94&39&1&0&0&0\\ 16&\!106&\!155&\!125&20&2&0&0\\ 17&54&\!132&\!184&89&15&0&0\\ 18&15&66&\!141&103&96&5&0\\ 19&0&16&63&63&114&50&0\\ 20,\!21&0&0&17&24&75&154&20\\ 22..28&0&0&0&0&0&23&72\end{matrix}\vskip-5.69046pt$
@@ 7 Combinatorics of the ML degree
Discussion of 7.1:.
We already conjectured that every copious graph is topologically copious. So, we assume the validity of that conjecture, and we consider graphs $G$ that are topologically copious. By Theorem 5.3, the ML degree $\mu(G)$ equals $(-1)^{n-3}$ times the Euler characteristic of the very affine variety $\mathcal{M}_{G}\subset(\mathbb{C}^{*})^{r}$. Hence 7.1 is a topological statement, saying that the signed Euler characteristic of $\mathcal{M}_{G}$ is equal to (34).

### v2

@@ 3 The Scattering Matroid
We work in the setting of (5), where $\mathcal{V}_{G}\subset\mathcal{K}_{G}\times\mathcal{M}_{0,n}$. Our hypothesis says that the map to $\mathcal{K}_{G}$ is dominant and finite-to-one. This means that the fiber over a generic point in $\mathcal{M}_{0,n}$ does not lie in a proper subspace of $\mathcal{K}_{G}$. In particular, that fiber doesn’t lie in a coordinate subspace $\{s_{ij}=0\}$, that is, no edge $ij$ of $G$ is a co-loop of the scattering matroid $\mathcal{S}(G)$. In symbols, this is the first equality of (17). The fact that the map $\mathcal{V}_{G}\to\mathcal{K}_{G}$ is of finite degree also implies that the codimension of the scattering correspondence in $\mathcal{K}_{G}\times\mathcal{M}_{0,n}$ equals $n-3={\rm dim}(\mathcal{M}_{0,n})$. Its codimension in $\mathbb{P}^{|G|-1}\times\mathcal{M}_{0,n}$ is therefore ${\rm codim}(\mathcal{K}_{G})+n-3=\kappa(G)+n-3$. The latter codimension is precisely the rank of $\mathcal{S}(G)$, thus (17) holds.
@@
A graph $G\!\subseteq\!\binom{[n]}{2}$ is called matroidally copious if both (17) and (18)=(19) are satisfied. In particular, the rank of the scattering matroid $\mathcal{S}(G)$ satisfies the upper bound
@@
Proposition 3.6 states that conditions (17) and (18)=(19) are equivalent to the map $\mathcal{V}_{G}\to\mathcal{K}_{G}$ being finite-to-one. Indeed, (17)-(18) implies that the projection to $\mathcal{K}_{G}$ is dominant. In particular, generic fibers are non-empty. Since, by (17), we have $\dim\mathcal{V}_{G}=\dim\mathcal{K}_{G}$, this map is also of finite degree. To show that the map to $\mathcal{M}_{0,n}$ is dominant, the fiber over a generic point $u\in\mathbb{P}^{n-1}$ should be non-empty. This follows from (17). The corresponding linear system should have a nonzero solution. Equivalently, $\mathrm{rank}(S_{G}(u))<\min(|G|,2n)$ for a generic $u\in\mathbb{P}^{n-1}$. The left-hand side is the rank of the scattering matroid $\mathcal{S}(G)$, and by Lemma 3.4 this rank is bounded by 2n32n-3. Thus, it is enough to have $\mathrm{rank}(\mathcal{S}(G))<|G|$. ∎
@@ 5 Partial Compactifications
A graph $G\subseteq\binom{[n]}{2}$ is topologically copious if the map (30) is an embedding. This property is independent of the choice of basis for the kinematic lattice $\mathcal{K}_{G}\,\cap\,\mathbb{Z}^{|G|}$.
@@ 6 Graphs with Few Vertices
Our experimental results give us considerable confidence to conjecture that condition 4 is equivalent to conditions 1,2,3 in Theorem 1.3 also for graphs $G$ without a universal vertex.
@@
$\small\begin{matrix}\hskip 18.49988pt|G|\backslash\mu(G)\!\!\!\!&1&2&3&4&\!5,\!6\!&\!\!7..9\!\!&\!\!10..24\!\!\\ \!\!12&\!10&2&0&0&0&0&0\\ \!\!13&\!14&8&1&0&0&0&0\\ \!\!14&\!11&13&5&2&0&0&0\\ \!\!15&5&9&5&6&3&0&0\\ \!\!16&0&4&2&5&6&0&0\\ \!\!17&0&0&0&3&2&4&0\\ \!\!18..21&0&0&0&0&1&2&6\end{matrix}\qquad\,\,\,\,\footnotesize\begin{matrix}\qquad|G|\backslash\mu(G)\!\!\!\!&1&2&3,\!4&5,\!6&\!\!7..10\!\!&\!\!11..20\!\!&\!\!21..120\!\!\\ 13&0&1&0&0&0&0&0\\ 14&52&29&3&0&0&0&0\\ 15&\!109&95&39&1&0&0&0\\ 16&\!104&\!156&\!126&20&2&0&0\\ 17&52&\!133&\!185&89&15&0&0\\ 18&15&65&\!142&103&96&5&0\\ 19&0&16&63&63&114&50&0\\ 20,\!21&0&0&17&24&75&154&20\\ 22..28&0&0&0&0&0&23&72\end{matrix}\vskip-5.69046pt$
@@ 7 Combinatorics of the ML degree
Discussion of Conjecture 7.1:.
We already conjectured that every copious graph is topologically copious. So, we assume the validity of that conjecture, and we consider graphs $G$ that are topologically copious. By Theorem 5.3, the ML degree $\mu(G)$ equals $(-1)^{n-3}$ times the Euler characteristic of the very affine variety $\mathcal{M}_{G}\subset(\mathbb{C}^{*})^{r}$. Hence Conjecture 7.1 is a topological statement, saying that the signed Euler characteristic of $\mathcal{M}_{G}$ is equal to (34).